# R2AG: Incorporating Retrieval Information into Retrieval Augmented Generation

### v1

@@ Abstract
Fuda Ye1, Shuangyin Li1,thanks: Corresponding author. The source code is available at https://github.com/yefd/RRAG.git., Yongqi Zhang2, Lei Chen2,3 1South China Normal University 2The Hong Kong University of Science and Technology (Guangzhou) 3The Hong Kong University of Science and Technology fudayip@m.scnu.edu.cn, shuangyinli@scnu.edu.cn, yongqizhang@hkust-gz.edu.cn, leichen@cse.ust.hk
@@ 2.2 Enhanced RAG
Notably, the above methods target providing higher-quality retrieval results to LLMs and actually treat retrieval and generation as two distinct processes. This separation fails to bridge the semantic gap between retrievers and LLMs fully.
@@ 3.2 Retrieval Feature Extraction
Before generation, it is necessary to obtain high-quality retrieval features. In R2AG, we first get semantic representations from the retriever $f_{R}$. Formally, a query $q$ and document $d$ are encoded into representations as $\mathbf{x}^{q}{=}f_{\mathbf{R}}(q)$ and $\mathbf{x}^{d}{=}f_{\mathbf{R}}(d)$, respectively. However, these representations can not be directly used because a single representation can not capture interactive features for LLM’s generation. Moreover, to suit various retrievers, it is intuitive to transform representations in different spaces into unified format features.
@@ 4.1 Datasets and Metrics
We evaluate R2AG on five datasets: Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), MuSiQue Trivedi et al. (2021), 2WikiMultiHopQA (2Wiki) Ho et al. (2020), and DuReader He et al. (2018). For NQ dataset, we utilize NQ-10, NQ-20, and NQ-30 datasets built by Liu et al. (2023), which contain 10, 20, and 30 total documents, respectively. DuReader is a multiple documents QA version built by Bai et al. (2023b). Detailed introduction and statistics are shown in Appendix D.
@@
Table 1: Main results on four English datasets. Results marked in grey background indicate the performance of foundation LLMs. Results in gray represent the performance of closed-source LLMs. Results in bold and results in underlined mean the best and second-best performance among current classified methods.
@@ 4.2 Baselines
The implementation details are in Appendix E.
@@ 4.4 Ablation Studies
Figure 5: Performance of R2AG7B and R2AG13B. Darker parts mean the difference value of R2AG13B.
@@ 4.5 Discussions
Further discussions and case studies are available in Appendix F and G.
@@ Appendix A Retrieval Feature Extraction Details
Datasets Language # Query # Train/Test # Tokens # Rel/Docs MAP
NQ-10 English 2655 2124/531 $\sim$2k 1/10 0.9602
NQ-20 English 2655 2124/531 $\sim$4k 1/20 0.9287
NQ-30 English 2655 2124/531 $\sim$6k 1/30 0.9215
HotpotQA English 97852 90447/7405 $\sim$2k 2.36/10 0.9138
MuSiQue English 22355 19938/2417 $\sim$3k 2.37/20 0.5726
2Wiki English 180030 167454/12576 $\sim$2k 2.42/10 0.9637
DuReader Chinese 200 160/40 $\sim$16k 1.82/20 0.7169
Table 4: Statistics of datasets. “# Rel/Docs” denotes the number of relevant documents and the total number of documents for each query. “MAP” represents the Mean Average Precision, a common retrieval metric.
@@ Appendix B Prompt Templates
Appendix C R2AG Algorithm
Algorithm 1 is the implementation of R2AG.
Input : Query $q$, retrieved documents $\mathcal{D}{=}\{d_{i}\}_{i=1}^{k}.$
Output : Answer $y$.
1Initialize R2-Former parameters $\Theta_{R}$, retriever $f_{\mathbf{R}}$, and LLM $f_{\mathbf{G}}$.
2Get query and documents semantic representation ($\mathbf{x}^{q}$ and $\mathbf{x}^{d}$) from retriever $f_{\mathbf{R}}$.
// Retrieval feature extraction
3 for $d_{i}$ in $D$ do
       Calculate $\text{input}_{i}$. // {Eq.10, 11, 12}
4
5 end for
6
// R2-Former
Calculate hidden states H; // {Eq. 2}
7
// Retrieval-aware prompting
Calculate hidden state $\textbf{E}^{R}$; // {Eq. 3}
8
Embedding lookup for text tokens; // {Eq. 4}
9
Fuse into final input embeddings E; // {Eq. 5}
10
Get the response ${y}$ from LLM $f_{\mathbf{G}}$; // {Eq. 6}
11
// Training strategy
12 if is training then
       Project the relevance scores $S$; // {Eq. 7}
13
      Calculate $\mathcal{L}_{QDM}$; // {Eq. 8}
14
15      Calculate $\mathcal{L}_{LM}$;
      Calculate final loss $\mathcal{L}$; // {Eq. 9}
16
17      Update parameters.
18 end if
19
return ${y}$
Algorithm 1 R2AG
Appendix D Dataset Introduction
@@
Detailed statistics can be found in Table 4.
@@
Appendix E Implementation Details
@@
Appendix F Further Discussion
Different Document Number
Figure 7 displays the performance comparison with different document numbers on NQ-30 dataset using LongChat1.57B. For both models, there is a notable trend where the performance metric generally decreases as the number of documents increases. However, this decline is steeper for LongChat1.57B compared to R2AG. It indicates that R2AG is particularly effective in scenarios where the number of documents can vary.
Figure 7: Performance comparison with different document numbers on NQ-30 dataset.
Efficiency of R2AG
In default settings, R2AG is designed to train only the R2-Former and the projection layer. The time complexity primarily depends on the number of the retrieval documents ($k$), the number of features (3 in our work), and the size of hidden layer in Transformer ($h_{1}$). In Table 6, we analyze the parameter count, including trainable and total parameters. The results indicate that the R2-Former only has few trainable parameters, and resource requirements of R2AG are substantially low. In the inference stage, R2AG incorporates retrieval information into LLMs’ embedding space, which theoretically increases the embeddings with $k$ tokens. To provide a practical comparison, we calculate the inference time on 2Wiki dataset using a single A800 GPU. The results indicate that the inference time increases by only 0.8%, making R2AG suitable for real-world applications. In practical scenarios, retrieved documents can be pre-encoded and stored in a vector database using tools like Faiss Douze et al. (2024) for efficient inference.
Methods # Trainable Params # Total Params Inference Time($k{=}10$)
LLaMA27B 0 6.738B 3.154s
R2-Former 1.319M 1.319M 0.499ms
R2AG 2.372M 6.741B 3.180s($\uparrow$ 0.8%)
Table 6: Efficiency of R2AG.
Appendix G Case Studies
Table 7 presents a case study on HotpotQA dataset, comparing R2AG with compression-based methods. In this case, the 1st and 3rd documents are relevant, and the 2nd document is a hard distractor. Only R2AG successfully provides the correct answer. In contrast, LongLLMLingua compresses the documents but retains their original order, which leads to ineffective outputs. Meanwhile, RECOMP confuses the 2nd and 3rd documents, resulting in hallucinations. More cases compared R2AG with standard RAG are shown in Table 8.
Case Question:Bordan Tkachuk was the CEO of a company that provides what sort of products? Relevant Documents: 1st, 3rd documents Gold Answer: IT products and services
R2AG Input Documents [1]similarity: <R>(Title: Bordan Tkachuk) Bordan Tkachuk ( ) is a British business executive, the former CEO of Viglen, also known from his appearances on the BBC-produced British version of "The Apprentice," interviewing for his boss Lord Sugar. [2]similarity: <R>(Title: BankUnited) BankUnited, Inc., with total consolidated assets of \$27.9 billion at December 31, 2016, is a bank holding company with one wholly owned subsidiary, BankUnited, collectively, the Company. BankUnited, a national banking association headquartered in Miami Lakes, Florida, provides a full range of banking services to individual and corporate customers through 94 banking centers located in 15 Florida counties and 6 banking centers in the New York metropolitan area. The Bank also provides certain commercial lending and deposit products on a national platform. The Company endeavors to provide, through experienced lending and relationship banking teams, personalized customer service and offers a full range of traditional banking products and services to both commercial and retail customers. [3]similarity: <R>(Title: Viglen) Viglen Ltd provides IT products and services, including storage systems, servers, workstations and data/voice communications equipment and services. [4]similarity: <R>(Title: Cardinal Health) Cardinal Health, Inc. is a Fortune 500 health care services company based in Dublin, Ohio. The company specializes in distribution of pharmaceuticals and medical products, serving more than 100,000 locations. The company also manufactures medical and surgical products, including gloves, surgical apparel and fluid management products. In addition, it operates the nation’s largest network of radiopharmacies. Cardinal Health provides medical products to over 75 percent of hospitals in the United States. In December 2013, it was announced that Cardinal Health would team up with CVS Caremark, which would form the largest generic drug sourcing operation in the United States. The venture was named Red Oak Sourcing and began operations in July 2014. [5]similarity: <R>(Title: Kingstone Companies) Kingstone Companies, Inc., which has its headquarters … [6]similarity: <R>(Title: Compass Minerals) Compass Minerals International, Inc is a United States … [7]similarity: <R>(Title: Jonathan Michael Ansell) Jonathan M. Ansell (born August 13, 1950 … [8]similarity: <R>(Title: Terren Peizer) Terren Scott Peizer dubbed the "Zelig of Wall Street" … [9]similarity: <R>(Title: Owens Minor) Owens & Minor () is a Fortune 500 company based in … [10]similarity: <R>(Title: Clean Power Finance) Clean Power Finance, headquartered in San Francisco … LLM’s Response IT products and services, including storage systems, servers, workstations and data/voice communications equipment and services.
LongLLMLingua Compressed Prompt (Title: Bordan Tkachuk) Bordan Tkachuk ( ) is a British business executive, the former CEO of Viglen, also known from his appearances on the BBC-produced British version of "The Apprentice," interviewing for his boss Lord Sugar. (Title: Cardinal Health) Cardinal Health, Inc. is a Fortune 500 health care services company based in Dublin, Ohio. The company specializes in distribution of pharmaceuticals and medical products, serving more than 100,000 locations. The company also manufactures medical and surgical products, including gloves, surgical apparel and fluid management products. In addition, it operates the nation’s largest network of radiopharmacies. Cardinal Health provides medical products to over 75 percent of hospitals in the United States. In December 2013, it was announced that Cardinal Health would team up with CVS Caremark, which would form the largest generic drug sourcing operation in the United States. The venture was named Red Oak Sourcing and began operations in July 2014. (Title Clean Power Finance) Clean Power Finance headquartered in San Francisco, California, is a financial services and software company for the residential solar industry. Clean Power Finance operates the CPF Market, an online business-to-business platform that connects institutional investors and lenders with residential solar professionals who need solar finance products to grow their businesses. Clean Power Finance provides the solar industry with CPF Tools, a solar sales, quoting and design software-as-a-Service (SaaS) solution: qualified residential solar channel partners access finance products through the software. Third-party investors create solar project finance funds; Clean Power Finance provides origination, underwriting and asset management services to the fund investors and markets investor capital to solar professionals as residential finance products, including solar leases and power purchase agreements (PPAs). (: Viglen) Viglen IT products and services, including storage systems, servers, workst and/voice communications and services. Title: Ow amp; Minor) Owens Minor () Fortune 500 company based Mechanicsville, Virginia, just of Richmond Owens Inc is health logistics specializing ining of healthcare products for hospitals Owens andt́agline is "Connect the World of Medical Product to Point of Care". The company supply chain services to healthcare providers and manufacturers healthcareable supplies as well devices and impl. Owens Minor logistics operations the United States Europe, where three quarters healthcare spending occurs. Ow Minor’ customers span the healthcare market independent hospitals large integrated healthcare networks as well group purchasing organizations, healthcare products manufacturers, the federal government. Owens & Minor is headquartered in Mechanicsville, Virginia and has annualized reven exceeding \$9 billion :stone Compies) Kingstone Companies,., which its headquarters in Kingston New York, property and casualty insurance … :Un) BankUn, Inc., with total consolidated assets \$2. billion at December 3 2016, is a bank holding wh owned subsidiary … : Compass Minerals) Compass Minerals International, Inc a United public company that its subsidiaries is leading producer … LLM’s Response No valid answer.
RECOMP Compressed Prompt Bordan Tkachuk is a British business executive. He was the CEO of Viglen. He was also known from his appearances on the BBC-produced British version of "The Apprentice," interviewing for his boss Lord Sugar. The company provides a variety of products and services. It offers IT products and services, including storage systems, servers, workstations, and data/voice communications equipment and services. The company also provides health care services, indicating a focus on the medical industry. The company manufactures medical and surgical products, including gloves, surgical apparel and fluid management products. … LLM’s Response Answer: health care services company
Table 7: Case studies of R2AG and compression-based methods.
Case Question:When did the animated series Kent Scott wrote end after beginning in September of 2002 on "Nick on CBS"? Relevant Documents: 1st, 4th documents Gold Answer: November
R2AG Input Documents [1]similarity: <R>(Title: Hawaii Five-0 (2010 TV series)) Hawaii Five-0 is an American action police procedural television series, which premiered on Monday, September 20, 2010 on CBS. The series is a reboot of the original series, which aired on CBS from 1968 to 1980. Like the original, it follows an elite state police task force set up to fight crime in the state of Hawaii. The series is produced by K/O Paper Products and 101st Street Television in association with CBS Productions, originally an in-name-only unit of but folded into CBS Television Studios, which has produced the series since the beginning of season three. On March 25, 2016, CBS renewed the series for a seventh season, which premiered on September 23, 2016. On March 23, 2017, CBS renewed the series for an eighth season, which premiered on September 29, 2017. [2]similarity: <R>(Title: Pelswick) Pelswick is an animated television series co-produced … [3]similarity: <R>(Title: Kenn Scott) Kenn Scott is a Toronto-based screenwriter noted for … [4]similarity: <R>(Title: Scott McCord) William Scott McCord (born April 19, 1971) is a Canadian actor, voice actor, musician, and composer currently based in Brooklyn. A versatile performer, he has appeared in the films "16 Blocks", "Shoot Ém Up", and the animated feature "The Nut Job". On television, he is best known for his voice over work in popular animated series. He plays Dan Kuso in "Bakugan Battle Brawlers", Tetsuya Watarigani in "Beyblade Metal Fusion", Owen and Trent in Fresh Animationś "Total Drama" series, Skull Boy in "Ruby Gloom", McGee in "Camp Lakebottom", and Jake in the animated TVOKids/Nickelodeon television series, "PAW Patrol". He is also Co-Artistic Director of Toronto-based Criminal Theatre, along with playwright and actor Rosa Laborde. In 2016, he won the Canada Screens Award for Best Performance in an Animated Series for his role as Squidgy on Guruś Justin Time. In 2002 he was nominated for a Dora Mavor Moore Award with Best Performance by a Male in Independent Theatre for his performance in Jacob Richmondś The Qualities of Zero. He has released two albums, the solo effort Blues For Sunshine (2009) and Scott McCord and the Bonafide Truth (2012). The latter is also the name of the soul/jazz/blues/rock eight piece band that was created for touring the first album. Scott McCord and the Bonafide Truth were nominated for Best Artist of the Year at the 2010 Maple Blues Awards. [5]similarity: <R>(Title: Project G.e.e.K.e.R.) Project G.e.e.K.e.R. is an animated television … [6]similarity: <R>(Title: Dora the Explorer) Dora the Explorer is an American educational … [7]similarity: <R>(Title: Duane Capizzi) Duane Capizzi is an American writer and television … [8]similarity: <R>(Title: Horseland (TV series)) Horseland is an American animated series … [9]similarity: <R>(Title: Henry Gilroy) Henry Gilroy is an American television screenwriter … [10]similarity: <R>(Title: David W. Zucker) David W. Zucker is President of Television … LLM’s Response November of that year.
LLaMA27B Input Documents As above but without retrieval information. LLM’s Response Pelswick
Case Question:When was the female co-founder of the organization that published the "Daily Word" devotional born? Relevant Documents: 2nd, 3rd documents Gold Answer: August 6, 1845
R2AG Input Documents [1]similarity: <R>(Title: The Word For Today) The Word For Today (known as The Word For You Today in some countries) … [2]similarity: <R>(Title: Unity Church) Unity, known informally as Unity Church, is a New Thought Christian organization that publishes the "Daily Word" devotional publication. It describes itself as a "positive, practical Christianity" which "teach[es] the effective daily application of the principles of Truth taught and exemplified by Jesus Christ" and promotes "a way of life that leads to health, prosperity, happiness, and peace of mind." [3]similarity: <R>(Title: Myrtle Fillmore) Mary Caroline "Myrtle" Page Fillmore (August 6, 1845 - October 6, 1931) was co-founder of Unity, a church within the New Thought Christian movement, along with her husband Charles Fillmore. Prior to that time, she worked as a schoolteacher. [4]similarity: <R>(Title: Daily Word) The Daily Word is a daily inspirational message provided … [5]similarity: <R>(Title: My Utmost for His Highest) My Utmost for His Highest is a daily devotional by … [6]similarity: <R>(Title: Our Daily Bread Ministries) Our Daily Bread Ministries (formerly RBC Ministries) is … [7]similarity: <R>(Title: Daily Light on the Daily Path) Daily Light on the Daily Path or Daily Light is … [8]similarity: <R>(Title: Our Daily Bread) Our Daily Bread is a devotional calendar-style booklet … [9]similarity: <R>(Title: The Clear Word) The Clear Word, originally published in March 1994 as … [10]similarity: <R>(Title: Weili Dai) Weili Dai () is a Chinese-born American businesswoman. She is the Director, … LLM’s Response August 6, 1845"
LLaMA27B Input Documents As above but without retrieval information. LLM’s Response The female co-founder of the organization that published the "Daily Word" devotional was born on 1952-09-20.
Case Question:What pop band features an Argentine actress that appeared in the television series "Chiquititas"? Relevant Documents: 1st, 4th documents Gold Answer: Erreway
R2AG Input Documents [1]similarity: <R>(Title: Luisana Lopilato) Luisana Lopilato ] (born May 18, 1987) is an Argentine actress and model. She has appeared in the television series "Chiquititas", "Rebelde Way", "Alma Pirata", "Casados con Hijos" and "Atracción x4". [2]similarity: <R>(Title: Celeste Cid) María Celeste Cid (] ; born 19 January 1984) is an … [3]similarity: <R>(Title: Nadia Di Cello) Nadia Mariel Di Cello (] ; born January 20, 1989 … [4]similarity: <R>(Title: Camila Bordonaba) Camila Bordonaba Roldán (] ; born September 4, 1984 in El Palomar, Buenos Aires, Argentina) (nicknamed Cami or Cato) is an Argentine actress, singer-songwriter, dancer, musician, Theatre director and former occasional model. She is best known for her roles in "Chiquititas", "Rebelde Way", "Son de Fierro" and "Atracción x4" and as a member of Argentine band Erreway. [5]similarity: <R>(Title: Milagros Flores) María de los Milagros Flores (born 26 April 1990 … [6]similarity: <R>(Title: Romina Yan) Romina Yankelevich (5 September 1974 – 28 September 2010) … [7]similarity: <R>(Title: Sweet Baby (Erreway song)) "Sweet Baby" is the debut single … [8]similarity: <R>(Title: Marcela Kloosterboer) Marcela Kloosterboer (born 5 July 1983) is … [9]similarity: <R>(Title: Mili Urién) Milagros Urién, simply known as Mili, is a fictional … [10]similarity: <R>(Title: Rincón de Luz (orphanage)) Rincón de Luz is a fictional orphanage and … LLM’s Response Erreway
LLaMA27B Input Documents As above but without retrieval information. LLM’s Response Luisana Lopilato
Table 8: Case studies of R2AG.

### v2

@@ Abstract
Fuda Ye1, Shuangyin Li1,thanks: Corresponding author. The source code is available at https://github.com/yefd/RRAG.git., Yongqi Zhang2, Lei Chen2,3 1School of Computer Science, South China Normal University 2The Hong Kong University of Science and Technology (Guangzhou) 3The Hong Kong University of Science and Technology fudayip@m.scnu.edu.cn, shuangyinli@scnu.edu.cn, yongqizhang@hkust-gz.edu.cn, leichen@cse.ust.hk
@@ 2.2 Enhanced RAG
Notably, the above methods target providing higher-quality retrieval results to LLMs and actually treat retrieval and generation as two distinct processes. This separation fails to bridge the semantic gap between retrievers and LLMs fully. Some approaches Deng et al. (2023); Sachan et al. (2021) enhance LLM comprehension abilities by incorporating documents into latent representations. However, these methods are typically designed for encoder-decoder LLMs, and constrain their suitability for prevailing decoder-only LLMs. While joint modeling methods Glass et al. (2022); Izacard et al. (2024) benefit from the joint optimization of LLMs and retrievers, they need extra training to align semantic spaces, which may hamper the generality of LLMs Zhao et al. (2024). Compared with these joint modeling methods, a key difference is that R2AG offers a cost-effective and non-destructive manner to bridge the semantic gap between LLMs and retrievers.
@@ 3.2 Retrieval Feature Extraction
Before generation, it is necessary to obtain high-quality retrieval features. In R2AG, we first get semantic representations from the retriever $f_{\mathbf{R}}$. Formally, a query $q$ and document $d$ are encoded into representations as $\mathbf{x}^{q}{=}f_{\mathbf{R}}(q)$ and $\mathbf{x}^{d}{=}f_{\mathbf{R}}(d)$, respectively. However, these representations can not be directly used because a single representation can not capture interactive features for LLM’s generation. Moreover, to suit various retrievers, it is intuitive to transform representations in different spaces into unified format features.
@@ 4.1 Datasets and Metrics
We evaluate R2AG on five datasets: Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), MuSiQue Trivedi et al. (2021), 2WikiMultiHopQA (2Wiki) Ho et al. (2020), and DuReader He et al. (2018). For NQ dataset, we utilize NQ-10, NQ-20, and NQ-30 datasets built by Liu et al. (2023), which contain 10, 20, and 30 total documents, respectively. DuReader is a multiple documents QA version built by Bai et al. (2023b). Detailed introduction and statistics are shown in Appendix C.
@@
Table 1: Main results on four English datasets. All enhanced RAG methods utilize the same foundation LLMs, with results marked in gray background indicating the performance of these foundation LLMs. Results in gray represent the performance of closed-source LLMs. Results in bold and results in underlined mean the best and second-best performance among current classified methods.
@@ 4.2 Baselines
The implementation details are in Appendix D.
@@ 4.3 Main Results
Figure 5: Performance of R2AG7B and R2AG13B. Darker parts mean the difference values of R2AG13B.
@@ 5 Conclusion and Future Work
Acknowledgments
This work was supported by Major Program of National Language Committee (WT145-39), Natural Science Foundation of Guangdong (2023A1515012073) and National Natural Science Foundation of China (No. 62006083).
@@ Appendix B Prompt Templates
Appendix C Dataset Introduction
@@
The ground truth labels are provided in original datasets. Detailed statistics can be found in Table 4.
@@
Appendix D Implementation Details